OpenStudy (anonymous):

The height of a parallelogram is 10 feet more than its base. If the area of the parallelogram is 120 square feet, find it's base and height. How would I set this up, and solve?

OpenStudy (anonymous):

area of a parallelogram = (base)(height) base = x height = x + 10 area = 120 x(x +10)=120

OpenStudy (anonymous):

x(x + 10) = 120 x(x + 10) -120 = 0 \[x ^{2} + 10x -120 = 0\] now us the quadratic equation \[(-b \pm \sqrt{b ^{2} - 4ac})/(2a)\]